OpenStudy (anonymous):

Compute the loss on shares of stock that are purchased at 27 3/8 and sold at 15 5/8.

OpenStudy (anonymous):

\[27+\frac{3}{8}-\left(15+\frac{5}{8}\right)=\frac{47}{4}=11\frac{3}{4} \]